OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 21 and the last term is 1,240,029?

OpenStudy (anonymous):

sum = a(1-r^n)/(1-r) http://en.wikipedia.org/wiki/Geometric_series. solve.

OpenStudy (anonymous):

????

OpenStudy (cwrw238):

in ronncc's formula a = first term = 21, n = 6 (number of terms), and r = common ratio

OpenStudy (anonymous):

is it 1527890?

OpenStudy (cwrw238):

the terms or of the form a, ar , ar^2, ar^3 ..... 6th term = ar^5 r^5 = 6th term / first term r^5 = 1,240,029 / 21 = 59049 r = 9 so plug these values into the formula: sum = 21 * 9^6 - 1 ------ = 6 - 1

OpenStudy (cwrw238):

= 2232048

OpenStudy (cwrw238):

where are you van??